# Quantum automorphism groups of 2-graphs

### v2

@@ 3. Main Results
Let $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ be two equivalent triples in the sense of Definition 2.7 so that for $i=1,2$, $\mathcal{G}_{i},\mathcal{H}_{i}$ have the common set of vertices $V$; vertex matrices of $\mathcal{G}_{i}$, $\mathcal{H}_{i}$ are $A_{i},B_{i}$ respectively. Then
${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})\cong{\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2}).$
In fact, ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ can be obtained from ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ by a similarity transformation.
Proof.
Let $T:V\rightarrow V$ be the bijection establishing the equivalence. Then $T_{\ast}:C(V)\rightarrow C(V)$ is an algebra isomorphism. We prove that if $\mathbb{G}$ acts faithfully on the triple $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, then it also acts faithfully on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. As usual we denote the vertex matrices of the graphs $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $A_{i}$, $B_{i}$ respectively; the edge sets of $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $E_{i}$, $F_{i}$ respectively. If $\alpha:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is the corresponding action then we claim that $\gamma:=(T_{\ast}^{-1}\otimes{\rm id})\circ\alpha\circ T_{\ast}:C(V)\rightarrow C(V)\otimes C(\mathbb{G})$ is an action on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Since $T_{\ast}$ is an algebra isomorphism, $\gamma$ is a $C^{\ast}$-action on the $C^{\ast}$-algebra $C(V)$. The corresponding magic unitary for the action $\gamma$ is $T^{-1}UT$ if $U$ is the magic unitary for the action $\alpha$ with respect to some basis of $C(V)$, where $T$ is the permutation matrix with respect to the same choice of basis. As $\mathbb{G}$ is a quantum symmetry of $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, we have $UA_{1}=A_{1}U$, $UB_{1}=B_{1}U$. As $TA_{2}=A_{1}T$, we have
$(T^{-1}UT)A_{2}=T^{-1}UA_{1}T=A_{2}(T^{-1}UT).$
Similarly using $TB_{2}=B_{1}T$ one can show that $(T^{-1}UT)B_{2}=B_{2}(T^{-1}UT)$. Hence $\gamma$ is a simultaneous quantum symmetry of the graphs $\mathcal{G}_{2},\mathcal{H}_{2}$. So we are left with proving the following commutative diagram:
{tikzcd}
where $\beta$ and $\beta^{\rm op}$ are restrictions of the corepresentation $\gamma^{\otimes 4}$. But it is easy to see that
$\gamma^{\otimes 4}=(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}.$
Then using the commutation $T^{\otimes 4}_{\ast}\circ\theta_{2\ast}=\theta_{1\ast}\circ T^{\otimes 4}_{\ast}$ and the fact that $\mathbb{G}$ is a quantum symmetry of $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$, we get
$\beta^{\rm op}\circ\theta_{2\ast}$ $=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}\circ\theta_{2\ast}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ\theta_{1\ast}\circ T_{\ast}^{\otimes 4}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ(\theta_{1\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(T^{-1^{\otimes 4}}_{\ast}\circ\theta_{1\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(\theta_{2\ast}\otimes{\rm id})\circ(T^{-1^{\otimes 4}}_{\ast}\otimes{\rm id})\circ\alpha^{\otimes 4}\circ T_{\ast}^{\otimes 4}$
$=$ $(\theta_{2\ast}\otimes{\rm id})\circ\beta.$
This proves that $\mathbb{G}$ acts on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Faithfulness is clear. So ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ acts on $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Therefore by the universal property, ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$. Similarly considering $T^{-1}$, one can show that ${\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ is a quantum subgroup of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ i.e.
${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})\cong{\rm Aut}^{+}(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2}).$
The assertion about the similarity transformation is clear. $\Box$
Now we are ready to define the notion of the quantum automorphism group of a $2$-graph.
@@
Let $\Lambda$ be a $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$. Then the quantum automorphism group of $\Lambda$ is defined to be the quantum automorphism group of $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ as defined in Definition 3.4.
Remark 3.8.
As explained earlier, thanks to Theorem 3.6 and Theorem A.1, the notion of the quantum automorphism group of a $2$-graph $\Lambda$ with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is well-defined. It will be denoted by either of the notations ${\rm Aut}^{+}(\Lambda)$ and ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$.
4. Computations
(a) Two copies of the same graph with the trivial bijection:
Recall the pullback of a $1$-graph $\mathcal{G}$ via the monoid morphism $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ given by $f(m,n)=m+n$. The pullback $f^{\ast}(\mathcal{G})$ is the $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, where $\mathcal{G}_{i}$’s are two copies of the same graph $\mathcal{G}$ and the bijection $\theta$ between the composable pairs is the trivial bijection. Since $\theta$ is the trivial bijection, replacing $(e^{2}_{\mu},e^{1}_{\nu})$ by $(e^{1}_{\alpha},e^{2}_{\beta})$ and $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ by $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ in the Equations (3.3), we see that (3.3) are redundant. Also since the individual $1$-graphs are two copies of the same graph, the following is immediate:
@@ 4. Computations
(b) Two different graphs with a nontrivial bijection:
@@
In fact the reader can verify that in this example there is a unique choice of $\theta$ given as above. The vertex matrices of the above two graphs $\mathcal{G}_{1},\mathcal{G}_{2}$ are respectively
$A_{1}=\begin{pmatrix}0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}$; $A_{2}=\begin{pmatrix}0&0&0&1\\ 0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\end{pmatrix}$.
With the usual notations, the action $\alpha$ of ${\rm Aut}^{+}(\mathcal{G}_{1}.\mathcal{G}_{2},\theta)$ on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji}$. Then the matrix $((q_{ij}))$ commutes with both the vertex matrices $A_{1},A_{2}$. A simple calculation reveals that the commutations impose the following relations among the generators $q_{ij}$:
@@
$=$ $q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}$
$=$ $q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
$=$ $q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
@@
where $q_{i}$’s are mutually orthogonal projections such that $\sum_{i=1}^{4}q_{i}=1$. Now it is easy to check that the above fundamental unitary corresponds to the CQG $C^{\ast}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$. Therefore ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is isomorphic to the classical group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
(c) Two copies of the graph $\mathcal{G}$ where $\mathcal{G}$ is the complete graph on four vertices:
{tikzpicture}
We denote one set of edges by $e_{i}$’s and the other set by $f_{i}$’s. The composable pairs are given by
$E\ \star\ E=\left\{\begin{matrix}&(e_{1},f_{2}),&(e_{1},f_{10}),&(e_{1},f_{7}),&(e_{2},f_{4}),&(e_{2},f_{12}),&(e_{2},f_{1}),\\ &(e_{3},f_{4}),&(e_{3},f_{12}),&(e_{3},f_{1}),&(e_{4},f_{3}),&(e_{4},f_{9}),&(e_{4},f_{5}),\\ &(e_{5},f_{6}),&(e_{5},f_{11}),&(e_{5},f_{8}),&(e_{6},f_{5}),&(e_{6},f_{3}),&(e_{6},f_{9}),\\ &(e_{7},f_{6}),&(e_{7},f_{11}),&(e_{7},f_{8}),&(e_{8},f_{2}),&(e_{8},f_{1}0),&(e_{8},f_{7}),\\ &(e_{9},f_{2}),&(e_{9},f_{10}),&(e_{9},f_{7}),&(e_{10},f_{3}),&(e_{10},f_{9}),&(e_{10},f5),\\ &(e_{11},f_{1}),&(e_{11},f_{4}),&(e_{11},f_{12}),&(e_{12},f_{6}),&(e_{12},f_{11}),&(e_{12},f_{8})\end{matrix}\right\}.$
@@
$(e_{1},f_{2})\mapsto(e_{4},f_{3}),$
$(e_{4},f_{3})\mapsto(e_{1},f_{2}),$
$(e_{8},f_{7})\mapsto(e_{6},f_{5}),$
$(e_{6},f_{5})\mapsto(e_{8},f_{7}),$
keeping rest of the elements of $E\star E$ fixed. As usual we write the action of ${\rm Aut}^{+}(\mathcal{G},\mathcal{G},\theta)$ on the set of vertices by
$\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji},\ {\rm for}\ i=1,\ldots,4.$
@@
For $(e,f)=(e_{1},f_{2})$ and $(f^{\prime},e^{\prime})=(e_{4},f_{3})$ the above equation reduces to
@@
We have the following equations for different choices of $(\mu,\nu)\in\{(e_{3},f_{4}),(e_{3},f_{1}),\\ (e_{3},f_{12}),(e_{2},f_{4}),(e_{2},f_{1}),(e_{2},f_{12}),(e_{11},f_{4}),(e_{11},f_{1}),(e_{11},f_{12})\}$:
$q_{21}q_{13}q_{21}=q_{21}q_{12}q_{21}$, $q_{21}q_{13}q_{31}=q_{21}q_{12}q_{31}$, $q_{21}q_{13}q_{41}=q_{21}q_{12}q_{41}$, $q_{31}q_{13}q_{21}=q_{31}q_{12}q_{21}$, $q_{31}q_{13}q_{31}=q_{31}q_{12}q_{31}$, $q_{31}q_{13}q_{41}=q_{31}q_{12}q_{41}$, $q_{41}q_{13}q_{21}=q_{41}q_{12}q_{21}$, $q_{41}q_{13}q_{31}=q_{41}q_{12}q_{31}$, $q_{41}q_{13}q_{41}=q_{41}q_{12}q_{41}$,
@@
which clearly shows that $q_{13}=q_{12}$. A similar calculation shows that $q_{43}=q_{42}$. Therefore the fundamental matrix takes the following form:
@@
Assume that $q_{11}=p$ and $q_{22}=q$, where $p$ and $q$ are projections. Then the magic condition forces $q_{41}=1-p=q_{14},q_{44}=p,q_{32}=1-q=q_{23},q_{33}=q$.
In the above calculation putting $(\mu,\nu)=(e_{8},f_{2})=(\nu^{\prime},\mu^{\prime})$ we see that

### v3

@@ 3. Main Results
Let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ and $(\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime},\theta^{\prime})$ be two equivalent triples in the sense of Definition 2.7. Then
${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)\cong{\rm Aut}^{+}(\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime},\theta^{\prime}).$
In fact, ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ can be obtained from ${\rm Aut}^{+}(\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime},\theta^{\prime})$ by a similarity transformation.
Now as the isomorphism class of a $2$-graph depends on the equivalence class of its defining triple (see the Appendix), the above theorem allows us to define the following.
@@
Let $\Lambda$ be a $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$. Then the quantum automorphism group of $\Lambda$ is defined to be the quantum automorphism group of $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ as defined in Definition 3.4. The quantum automorphism group will be denoted by either of the notations ${\rm Aut}^{+}(\Lambda)$ or ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$.
4. Examples and Computations I
(a) Two copies of the same graph with trivial bijection
Recall the pullback of a $1$-graph $\mathcal{G}$ via the monoid morphism $f:\mathbb{N}^{2}\rightarrow\mathbb{N}$ given by $f(m,n)=m+n$. The pullback $f^{\ast}(\mathcal{G})$ is the $2$-graph with the defining triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, where $\mathcal{G}_{l}$’s are two copies of the same graph $\mathcal{G}$ and the bijection $\theta$ between the composable pairs is the trivial bijection. Since $\theta$ is the trivial bijection, replacing $(e^{2}_{\mu},e^{1}_{\nu})$ by $(e^{1}_{\alpha},e^{2}_{\beta})$ and $(e^{2}_{\mu^{\prime}},e^{1}_{\nu^{\prime}})$ by $(e^{1}_{\alpha^{\prime}},e^{2}_{\beta^{\prime}})$ in (3.3), we see that (3.3) are redundant. Also since the individual $1$-graphs are two copies of the same graph, the following is immediate:
@@ 4. Examples and Computations I
(b) Two different graphs with a nontrivial bijection
@@
In fact, the reader can verify that in this example there is a unique choice of $\theta$ given as above. With the usual notations, the action $\alpha$ of ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ on $C(V)$ is given by $\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji}$. A simple calculation reveals that the commutation of the unitary $((q_{ij}))$ with the vertex matrices impose the following relations among the generators $q_{ij}$’s:
@@
$\quad=q_{t(e^{2}_{\mu^{\prime}})t(e^{2}_{\mu})}q_{s(e^{1}_{\nu^{\prime}})s(e^{1}_{\nu})}$
$\quad=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}$
$\quad=q_{t(e^{1}_{\alpha^{\prime}})t(e^{1}_{\alpha})}q_{s(e^{1}_{\alpha^{\prime}})s(e^{1}_{\alpha})}q_{s(e^{2}_{\beta^{\prime}})s(e^{2}_{\beta})}.$
@@
where $q_{i}$’s are mutually orthogonal projections such that $\sum_{i=1}^{4}q_{i}=1$. Now it is easy to check that the above fundamental unitary corresponds to the CQG $C^{\ast}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$. Therefore, ${\rm Aut}^{+}(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ is isomorphic to the classical group $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
(c) Two copies of the graph $\mathcal{G}$ where $\mathcal{G}$ is the complete graph on four vertices
[…]
The composable pairs are given by
$E\ \star\ E=\left\{\begin{matrix}&(e_{1},e_{2}),&(e_{1},e_{10}),&(e_{1},e_{7}),&(e_{2},e_{4}),&(e_{2},e_{12}),&(e_{2},e_{1}),\\ &(e_{3},e_{4}),&(e_{3},e_{12}),&(e_{3},e_{1}),&(e_{4},e_{3}),&(e_{4},e_{9}),&(e_{4},e_{5}),\\ &(e_{5},e_{6}),&(e_{5},e_{11}),&(e_{5},e_{8}),&(e_{6},e_{5}),&(e_{6},e_{3}),&(e_{6},e_{9}),\\ &(e_{7},e_{6}),&(e_{7},e_{11}),&(e_{7},e_{8}),&(e_{8},e_{2}),&(e_{8},e_{1}0),&(e_{8},e_{7}),\\ &(e_{9},e_{2}),&(e_{9},e_{10}),&(e_{9},e_{7}),&(e_{10},e_{3}),&(e_{10},e_{9}),&(e_{10},f5),\\ &(e_{11},e_{1}),&(e_{11},e_{4}),&(e_{11},e_{12}),&(e_{12},e_{6}),&(e_{12},e_{11}),&(e_{12},e_{8})\end{matrix}\right\}.$
@@
$(e_{1},e_{2})\mapsto(e_{4},e_{3}),$
$(e_{4},e_{3})\mapsto(e_{1},e_{2}),$
$(e_{8},e_{7})\mapsto(e_{6},e_{5}),$
$(e_{6},e_{5})\mapsto(e_{8},e_{7}),$
keeping rest of the elements of $E\star E$ fixed. As usual, we write the action of ${\rm Aut}^{+}(\mathcal{G},\mathcal{G},\theta)$ on the set of vertices by
$\alpha(\delta_{i})=\sum_{j=1}^{4}\delta_{j}\otimes q_{ji},\text{ for }i=1,\ldots,4.$
@@
For $(e,f)=(e_{1},e_{2})$ and $(f^{\prime},e^{\prime})=(e_{4},e_{3})$ the above equation reduces to
@@
We have the following equations for different choices of $(\mu,\nu)\in\{(e_{3},e_{4}),(e_{3},e_{1}),\\ (e_{3},e_{12}),(e_{2},e_{4}),(e_{2},e_{1}),(e_{2},e_{12}),(e_{11},e_{4}),(e_{11},e_{1}),(e_{11},e_{12})\}$:
$\begin{matrix}q_{21}q_{13}q_{21}=q_{21}q_{12}q_{21},&q_{21}q_{13}q_{31}=q_{21}q_{12}q_{31},&q_{21}q_{13}q_{41}=q_{21}q_{12}q_{41},\\ q_{31}q_{13}q_{21}=q_{31}q_{12}q_{21},&q_{31}q_{13}q_{31}=q_{31}q_{12}q_{31},&q_{31}q_{13}q_{41}=q_{31}q_{12}q_{41},\\ q_{41}q_{13}q_{21}=q_{41}q_{12}q_{21},&q_{41}q_{13}q_{31}=q_{41}q_{12}q_{31},&q_{41}q_{13}q_{41}=q_{41}q_{12}q_{41},\end{matrix}$
@@
which clearly shows that $q_{13}=q_{12}$. A similar calculation shows that $q_{43}=q_{42}$. Therefore, the fundamental matrix takes the following form:
@@
Assume that $q_{11}=p$ and $q_{22}=q$, where $p$ and $q$ are projections. Then the magic condition forces $q_{41}=1-p=q_{14},q_{44}=p,q_{32}=1-q=q_{23},q_{33}=q$. In the above calculation putting $(\mu,\nu)=(e_{8},f_{2})=(\nu^{\prime},\mu^{\prime})$ we see that